How many total atoms are in 0.650 g of P2O5?
1
Expert's answer
2019-02-21T03:23:05-0500

N=NA*n

n=m/M

M(P2O5) = 141.94 g/mol

n(P2O5) = 0.650 g/141.94g/mol = 0.00458 mol

N(P2O5) = 0.00458 mol * 6.022*1023 mol-1 = 2.76*1021 molecules

1 molecule of P2O5 contains 2+5 = 7 atoms

2.76*1021 molecules of P2O5 contain x atoms

1/2.76*1021 = 7/x

x= 1.93*1022 atoms

Answer: 1.93*1022